OpenStudy (anonymous):

solve 50 squared+ 60 squared =N . divided by 625

OpenStudy (maheshmeghwal9):

N=(2500+3600) x 625. Now use calculator.

OpenStudy (maheshmeghwal9):

or u find urself:)

OpenStudy (anonymous):

so 3812500

OpenStudy (maheshmeghwal9):

wait a minute:)

OpenStudy (anonymous):

?

OpenStudy (maheshmeghwal9):

ya:)

OpenStudy (maheshmeghwal9):

well done:-)

OpenStudy (maheshmeghwal9):

@theatreliver

OpenStudy (anonymous):

Divided, not multiplied. So N=(2500+3600) / 625.

OpenStudy (maheshmeghwal9):

Hmm....Plz see question carefully:)

OpenStudy (anonymous):

It says divided.

OpenStudy (maheshmeghwal9):

no divide is for R.H.S, when going to L.H.S. It becomes multiplied

OpenStudy (anonymous):

so it wold be 9.76?

OpenStudy (anonymous):

Yep :)

OpenStudy (maheshmeghwal9):

nonono.

OpenStudy (maheshmeghwal9):

wait a minute we should ask other person @ajprincess Plz help:)

OpenStudy (anonymous):

solve 50 squared+ 60 squared =N . divided by 625 ______

OpenStudy (maheshmeghwal9):

Am i wrong? @ajprincess

OpenStudy (maheshmeghwal9):

may be @thechocoluver445 right:)

OpenStudy (anonymous):

@theatreliver how is the problem written? i think the wording may be confusing everyone :)

OpenStudy (maheshmeghwal9):

ya @thechocoluver445 is right:)

OpenStudy (ajprincess):

I thnk u r right. coz that's vat I too thought. bt there is only one person who can say it correct or wrong. it's none other than @theatreliver.

OpenStudy (ajprincess):

*whether it is correct or wrong.

OpenStudy (anonymous):

50^2 + 60^2 =N / 625

OpenStudy (anonymous):

oh. then @maheshmeghwal9 is right. you would multiply by 625. 3,812,500 is the correct answer.

OpenStudy (maheshmeghwal9):

but @thechocoluver445 also thought very perfect. <My thought was only limited.

OpenStudy (this):

7.092

OpenStudy (anonymous):

7.092?

OpenStudy (anonymous):

@this elaborate a bit?

OpenStudy (maheshmeghwal9):

How @this ?

OpenStudy (maheshmeghwal9):

@theatreliver the answer told by myself is 100% correct:) u don't worry:)

OpenStudy (maheshmeghwal9):

Let us see what @this has to say?

OpenStudy (anonymous):

For some reason this's username cracks me up. Whenever she/he is typing a reply, it says 'this is typing a reply'. :3

OpenStudy (anonymous):

*she

OpenStudy (maheshmeghwal9):

?

OpenStudy (anonymous):

so im just going to go with 3,812,500 thanks guys (:

OpenStudy (maheshmeghwal9):

yeah very good:) u r a good mathsolver:)

OpenStudy (maheshmeghwal9):

sorry but I can't iunderstand what @this did?

OpenStudy (maheshmeghwal9):

actually*

OpenStudy (ajprincess):

she thought it as sqrt50. i suppose so.

OpenStudy (maheshmeghwal9):

k! no worries @this good try:)

OpenStudy (anonymous):

oh, i see what @this meant now.

OpenStudy (maheshmeghwal9):

According to u, u were right:)

OpenStudy (maheshmeghwal9):

@this

OpenStudy (this):

oops srry guys D:

OpenStudy (maheshmeghwal9):

no worries:-)

OpenStudy (ajprincess):

it's k. that happens to evryone. no worries.

OpenStudy (maheshmeghwal9):

so i think @theatreliver must be satisfied now:)

OpenStudy (anonymous):

what degree fahrenheit is it if the temperature is 15 degrees celsius? is this 59?

OpenStudy (maheshmeghwal9):

There is a formula but I don't remember that u should better search google for that:)

OpenStudy (anonymous):

F= C * 9 / 5 + 32 = 15 * 9 / 5 + 32 = 59

OpenStudy (maheshmeghwal9):

or wait for @thechocoluver445

OpenStudy (anonymous):

That's right.

OpenStudy (maheshmeghwal9):

gr8

OpenStudy (anonymous):

thanks (:

OpenStudy (maheshmeghwal9):

u deserve it:)